OpenStudy (anonymous):

What is f(x + h) when f(x) = 3x^2 + 2x - 4 ?

OpenStudy (amistre64):

replace each "x" in the original statement by "x+h"

OpenStudy (amistre64):

f(x+h)=3(x+h)^2+2(x+h)-4, expanded it out as wanted

OpenStudy (anonymous):

i did and got f(x + h) = 3x^2 + 3h^2 + 6xh + 2x + 2h - 4 this is just to make sure i am correct

OpenStudy (amistre64):

looks good to me

OpenStudy (anonymous):

thanks